1) A horizontal curve is being designed around a pond with a tangent length of 1200 ft and central angle of 0.5211 radians. If the PI is at station 145+00, determine the station of PT.
⦁   168+45.43
⦁   156+45.43
⦁   173+94.00
⦁   156+72.72

2) A car is traveling downhill on a suburban road with a grade of 3% at a speed of 35 mi/h. Determine the stopping sight distance based on a reaction time of 2.5 second, and a deceleration rate of 11.2 ft/s2.
⦁   257.08 ft
⦁   236.63 ft
⦁   249.15 ft
⦁   254.23 ft

3) A 400-ft equal-tangent sag vertical curve has its PVC at station 100+00 and elevation 500 ft. The initial grade is -4.0% and the final grade is +2.5%. Determine the elevation of the lowest point of the curve.
⦁   495.077 ft
⦁   495.250 ft
⦁   485.231 ft
⦁   492.043 ft

4) Which of the following is NOT a safety performance measure?

A. Crash rates
B. Number of traffic conflicts
C. Total traffic volume
D. Fatalities per accident
Answer #1

0 Given, T = 1200 ft I = 0.5211 radians 0.5211 x. 180 = 29.857 PI station = 145 +00 Radius of curve (R) (R.) = TL tan ( 2 ) NNow Chas of PT of Pl - Tu + L Chainage of hainage of 1 4500 - 1200+ 23 45. 443 156 45.4 43. N of PT = 156+ 45.44 3 Ars (2nd o3 Griven, L = 400th. pre stations 100too PVC elevation (pvc) = 500ft. g, 4 % 2 +2.5% I low = ? Distance of low point from pvc4 safety Total traffic volume is not a performance Thus, option (6) is the correct answer. ince measure.